# Human Stress Response and Perceived Safety during Encounters with Quadruped Robots

### v1

@@ I Introduction
There exist well known qualitative methods for evaluating stress and perceived safety [bartneck2009measurement]. Subjective questionnaires have been popularly employed in literature [hauser2023s, van2023increasing, taylor2022observer] to measure preference and human mental state during HRI. However, questionnaires have poor temporal resolution and are subject to many sources of bias including moderacy response bias and memory effects induced by repetitive administration. Questionnaires also impact HRI, requiring researchers to adopt rapidly administered questionnaires which have limited detail and coverage or to limit the choice of robots to be studied to those with a user interface for questionnaires. To overcome these limitations, more objective measures have also been considered in HRI [dehais2011physiological, chen2011touched, stark2018personal, willemse2017affective]. Non-invasive physiological sensing enables tracking of the human internal state in real-time, for example to evaluate stress during HRI [akalin2022you].
We hypothesize these sensors allow tracking perceived safety during human-robot encounters by decoding acute stress levels from known biomarkers in ECG and EDA data. This is significant in two ways: (1) To the best of our knowledge, this is the first analysis of physiological response to mobile robots navigating freely [rubagotti2022perceived]. Our study deploys two quadruped robots in a realistic apartment setting (Fig. 1). (2) We characterize human-robot encounters. This subset of human-robot interactions involve humans and robots coexisting in a shared space with awareness of one another, who are not actively interacting. Human-robot encounters represent situations that are likely to dominate total interaction time if mobile robots are deployed in home or other indoor environments. Our results provide a few insights, based on biological signal analysis, regarding the elements of human-robot encounters that affect the perceived safety of the robots by humans. These insights can be used in future design where perceived safety is a concern and pave the way for an online pipeline whereby the signals from a user’s sensors could be decoded to inform robot behavior in real-time [staffa2022enhancing, rani2004anxiety].
@@ II-A Stress Response and Comfort in HRI
Perceived safety during encounters with mobile robots is understudied [hauser2023s, akalin2023taxonomy]. Within existing work we identify three broad categories of studies concerning comfort and stress during HRI. The first group involves studying participant stress response using physiological measures. These studies integrate robots in human activities with known external stressors like seminar style presentations or cognitive quizzes to evoke discomfort [akalin2022you, staffa2022enhancing, turner2019use]. They ask how replacing humans with robots impacts people’s response to situations, however none consider mobile robots moving freely through shared environments. The second group of studies include more subjective measures, like the Godspeed questionnaire or Perceived Stress Scale [bartneck2009measurement, hauser2023s, koay2005methodological, cohen1994perceived]. The final group of studies particularly focus on specially trained workers, who collaborate with industrial robots on a regular basis [rubagotti2022perceived, villani2018survey]. Ultimately, the proposed study intersects the first two categories, using a combination of objective biological signals with subjective responses to questionnaires and interviews.
@@ II-B Stress Decoding from Physiological Signals
Non-invasive physiological recordings have been shown to contain a wealth of information about human physical and mental states. We are interested in mental stress, particularly acute stress from short recordings. This is fulfilled by electrodermal activity (EDA) recordings, or galvanic skin response, which measures changes in skin conductance over time on the palm, foot, or wrist. Skin conductance time series consists of two components: tonic and phasic. Tonic refers to the low-frequency changes in the baseline of the signal, while the phasic refers to the sharp spikes followed by recovery to baseline in the signal. The phasic component is known to correspond to psychological arousal events, allowing quantitative determination of mental stress over time. Decoding acute stress from tonic-phasic decompositions of EDA signals has been demonstrated extensively [greco2021acute].
Another signal to detect mental stress is heart rate variability (HRV) [kim2018stress], which can be evaluated from any modality for reliable heart beat information. HRV features have been linked to a variety of both acute and chronic stress in both real life and experimentally induced situations [thehrvreview]. This is due to the fact that HRV reflects complex changes in sympathetic and parasympathetic nervous systems which manage the homeostasis of psychological arousal. HRV is often evaluated over long recordings; however it can be reliably estimated from ultra-short-term and short-term recordings like our study [USTHRV1, USTHRV2]. Acute stress has been associated with decreases in HRV entropy in both experimentally induced situations [USTHRV2, brugnera_entropy] and real-life situations [dmitri_entropy, exam_entropy]. Decreases in SDNN and RMSSD and increases in low-frequency bandpower (0.04–0.15 Hz) compared to high-frequency (0.15–0.40 Hz) are also commonly noted features of acute stress [thehrvreview].
@@ III-C Experiment Procedure
First, participants are presented with a consent form. Upon signing consent, the chest e-tattoo and E4 wristwatch are applied to participants. Participants are seated together on the living room couch to fill out a pre-experiment questionnaire. The participants then watch a 5 minute relaxing nature video to establish a relaxed baseline state for physiological data. After, the participants’ seating locations are updated to isolated if necessary (chosen at random). Next, experimental trials begin, each characterized by (1) a seating position, (2) single or multi-robot and (3) search or navigation behavior. Between trials participants take a 10 question Likert scale questionnaire. It is a combination of surveys focused on stress, comfort and perceived safety used in [akalin2022you]. After four trials, participant seating is switched. The same 4 behaviors are again repeated in the new seating locations with order again randomized. Participants then watch the same 5 minute nature video, establishing a second known baseline relaxation state to confirm that perturbations in the physiological signals during the 8 trials were induced by the robot and not a confounding factor such as elapsed time in the experiment and fatigue. After the nature video, participants removed the sensors and were debriefed then interviewed.
@@ IV-B Exploratory Data Analysis
Next, we used the Chi-squared test to examine the 20 most statically significant features for the Baseline v. Robot condition. We note that EDA and HRV features were shortlisted in an approximately even proportion (9 EDA, 11 HRV) by the algorithm despite more candidate HRV features evaluated. This suggests that the multimodal approach based on both cardiac and electrodermal activity was informative for the estimation of robot-induced stress. Since exposure to robots is not a well-established stressor condition (in contrast to established psychological or neuroscientific tasks such as Stroop and Speech tests), an important step is to: (1) further examine the shortlisted features based on whether they have reasonable physiological background for being predictive of acute stress and arousal, and (2) examine the evolution of these features in the encounter conditions of interest. These steps will help to ensure that the predictive performance of the automatic human-robot encounter classifier in Section IV-C is not a result of overfitting.
@@ V Results
Fig. 7 demonstrates screenshots from several encounters with human participants. For data from all subjects combined, the Mann-Whitney-Wilcoxon two-sided test was used with Bonferroni correction for multiple comparisons. On average, we found that subjects were: more nervous whenever the robots were in the shared environment compared to when they were not; more nervous when the robots were travelling together rather than alone; and more nervous when the robots were navigating compared to searching (Fig. 6(a-c)). Whether the subjects were located alone or together did not induce significant changes in the stress features examined (Fig. 6(d)).
@@ VI Discussion
Fig. 6 and Tables LABEL:table:cm1 - LABEL:table:cm4 suggest that for the average human participant, sharing indoor spaces with mobile quadrupedal robots was associated with statistically significant physiological changes indicative of acute stress. Further, it was possible to distinguish, with higher than chance-level accuracy, specific characteristics (single v. multi robot and navigation v. search) about the human-robot encounters based on the ECG and EDA signals. These observations should be critically examined for potential insights for designers of robots and human-robot interaction.
For instance, we found that the average participant was more stressed by robots on navigation behavior than those on search behavior. One could hypothesize that this was due to robots on search inherently spending more time near the participants, giving them an opportunity to become more familiar with the robot. Indeed, familiarity, which is influenced by interaction duration and frequency, is one of the major factors behind perceived safety [akalin2023taxonomy]. However, to confirm this idea, further analysis of the data may be necessary (e.g. correlate human-robot distance with stress features to decouple navigation/search behavior from distance; interview and questionnaire data collected may contain information about how differently the subjects perceived robots on navigation vs. search). While less statistically significant, we also found that the average participant was more stressed encountering multiple robots travelling together rather than alone. This could suggest that whenever possible, mobile robots should travel alone, at least in similar physical settings as our experiment (i.e. closed indoors space occupied by humans), to maximize perceived safety. We also found that whether the subjects were alone or together was not associated with significant differences in physiological features of stress. This could indicate that mobile robots may not pose significant perceived safety issues even when deployed in environments where people work alone.
@@
VII Acknowledgments

### v2

@@ I Introduction
Subjective questionnaires have been popularly employed in literature to measure preceived safety, preference and human mental state during HRI [bartneck2009measurement, hauser2023s, van2023increasing, taylor2022observer]. However, questionnaires have poor temporal resolution and are subject to many sources of bias including moderacy response bias and memory effects induced by repetitive administration [choi2005peer]. Questionnaires also impact HRI, requiring researchers to adopt rapidly administered questionnaires which have limited detail and coverage or, in some cases, limit the choice of robots to be studied to those like the Pepper robot with a user interface for questionnaires [akalin2023taxonomy, rubagotti2022perceived]. To overcome these limitations, physiological measures have also been considered in HRI [dehais2011physiological, chen2011touched, stark2018personal, willemse2017affective]. Non-invasive physiological sensing enables the tracking of the human internal state in real-time, for example to evaluate stress during HRI [akalin2022you].
We hypothesize these sensors allow tracking perceived safety during human-robot encounters by decoding acute stress levels from known biomarkers in ECG and EDA data. This is significant in two ways: (1) To the best of our knowledge, this is the first analysis of physiological response to mobile robots navigating freely [rubagotti2022perceived]. Our study deploys two quadruped robots in a realistic apartment setting (Fig. 1). (2) We characterize human-robot encounters. This subset of human-robot interactions involve humans and robots coexisting in a shared space with awareness of one another, who are not actively interacting. Human-robot encounters represent situations that are likely to dominate total interaction time if mobile robots are deployed in home or other indoor environments. Our results provide a few insights, based on physiological signal analysis, regarding the elements of human-robot encounters that affect the perceived safety of the robots by humans. These insights can be used in future design where perceived safety is a concern and pave the way for an online pipeline whereby the signals from a user’s sensors could be decoded to inform robot behavior in real-time [staffa2022enhancing, rani2004anxiety].
@@ II-A Stress Response and Comfort in HRI
Perceived safety during encounters with mobile robots remains understudied in many domains [hauser2023s, akalin2023taxonomy]. The first group of works we find leverage more objective physiological measures to identify stress response. These studies integrate robots in human activities with known external stressors like seminar style presentations or cognitive quizzes to evoke discomfort [akalin2022you, staffa2022enhancing, turner2019use]. They ask how replacing humans with robots impacts people’s response to situations, however none consider mobile robots moving freely through shared environments. The second group of studies leverage more subjective measures, for example statistical analysis of subjective questionnaires [hauser2023s, van2023increasing, aliasghari2023we, wang2023effects, cohen1994perceived, mavrogiannis2022social] or qualitative evaluation of open form responses [doring2019love, fortunati2021rise, joffe2011thematic, terry2017thematic]. Ultimately, the proposed study differentiates from many works, using a combination of methods to investigate perceived safety during mobile robot encounters.
@@ II-B Stress Decoding from Physiological Signals
Non-invasive physiological recordings have been shown to contain a wealth of information about human physical and mental states [greco2021acute, thehrvreview]. We are interested in mental stress, particularly acute stress from short recordings. This is fulfilled by electrodermal activity (EDA) recordings, or galvanic skin response, which measures changes in skin conductance over time on the palm, foot, or wrist [zangroniz2017electrodermal, critchley2002electrodermal]. Skin conductance time series consists of two components: tonic and phasic. Tonic refers to the low-frequency changes in the baseline of the signal, while the phasic refers to the sharp spikes followed by recovery to baseline in the signal [greco2015cvxeda]. The phasic component is known to correspond to psychological arousal events, allowing quantitative determination of mental stress over time. Decoding acute stress from tonic-phasic decompositions of EDA signals has been demonstrated extensively [greco2021acute, zangroniz2017electrodermal, critchley2002electrodermal, caruelle2019use].
Another signal to detect mental stress is heart rate variability (HRV) [kim2018stress], which can be evaluated from any modality for reliable heart beat information. HRV features have been linked to a variety of both acute and chronic stress in both real life and experimentally induced situations [thehrvreview]. This is due to the fact that HRV reflects complex changes in sympathetic and parasympathetic nervous systems which manage the homeostasis of psychological arousal. HRV is often evaluated over long recordings; however it can be reliably estimated from ultra-short-term and short-term recordings like our study [USTHRV1, USTHRV2]. Acute stress has been associated with decreases in HRV entropy in both experimentally induced situations [USTHRV2, brugnera_entropy] and real-life situations [dmitri_entropy, exam_entropy]. SDNN (Standard Deviation of NN intervals) and RMSSD (Root Mean Square of Successive Differences) are both measures commonly used in the analysis of heart rate variability (HRV), which is the variation in the time intervals between successive heartbeats. Decreases in SDNN and RMSSD and increases in low-frequency bandpower (0.04–0.15 Hz) compared to high-frequency (0.15–0.40 Hz) are also commonly noted features of acute stress [thehrvreview].
@@ III-C Experiment Procedure
First, participants are presented with a consent form. Upon signing consent, the chest e-tattoo and E4 wristwatch are applied to participants. Participants are seated together on the living room couch to fill out a pre-experiment questionnaire. The participants then watch a 5 minute relaxing nature video to establish a relaxed baseline state for physiological data. After, the participants’ seating locations are updated to isolated if necessary (chosen at random). Next, experimental trials begin, each characterized by (1) a seating position, (2) single or multi-robot and (3) search or navigation behavior. Order effects of the conditions are addressed by counterbalancing. In particular, we randomize the isolated v. participant seating and each of the four robot behaviors within each seating condition. Between trials participants take a 10 question Likert scale questionnaire. It is a combination of surveys focused on stress, comfort and perceived safety used in [akalin2022you]. After four trials, participant seating is switched. The same 4 behaviors are again repeated in the new seating locations with order again randomized. Participants then watch the same 5 minute nature video, establishing a second known baseline relaxation state to confirm that perturbations in the physiological signals during the 8 trials were induced by the robot and not a confounding factor such as elapsed time in the experiment and fatigue. After the nature video, participants removed the sensors and were debriefed then interviewed.
@@ IV-B Exploratory Data Analysis
Next, we used the Chi-squared test to examine the 20 most statically significant features for the Baseline v. Robot condition. We chose a Chi-squared test compared with a principal component analysis (PCA) for interpretability of the results. While a PCA can be performed for a similar purpose, the resulting components cannot be compared with features of acute stress identified from literature as done in this work, which results in less certainty about the applicability of the results. This is due to the fact that there is no common “stress metric” in literature which can be used. In the feature selection via chi-squared test function available in Matlab, which is good to use for continuous data for classification problems, the features are ranked according to negative logs of the p-values. We note that EDA and HRV features were shortlisted in an approximately even proportion (9 EDA, 11 HRV) by the algorithm despite more candidate HRV features evaluated. This suggests that the multimodal approach based on both cardiac and electrodermal activity was informative for the estimation of robot-induced stress. Since exposure to robots is not a well-established stressor condition (in contrast to established psychological or neuroscientific tasks such as Stroop and Speech tests), an important step is to: (1) further examine the shortlisted features based on whether they have reasonable physiological background for being predictive of acute stress and arousal, and (2) examine the evolution of these features in the encounter conditions of interest. These steps will help to ensure that the predictive performance of the automatic human-robot encounter classifier in Section IV-C is not a result of overfitting.
@@ V Results
All of the data taken during experiments including from robots, microphones, and physiological sensors is published for open access at [dataset] and a video of experimental encounters can be found at https://youtu.be/3xu5AoBYi44. Quantitative results are also reported in the data repository. Fig. 7 demonstrates screenshots from several encounters with human participants. For data from all subjects combined, the Mann-Whitney-Wilcoxon two-sided test was used with Bonferroni correction for multiple comparisons. On average, we found that subjects were: more stressed whenever the robots were in the shared environment compared to when they were not; more stressed when the robots were travelling together rather than alone; and more stressed when the robots were navigating compared to searching (Fig. 6(a-c)). Whether the subjects were located alone or together did not induce significant changes in the stress features examined (Fig. 6(d)).
@@
In Fig 6, the values are adjusted with Bonferroni correction. The two-sided Mann-Whitney-Wilcoxon tests with Bonferroni correction that yields pairwise across-subjects significance is different from the p-values reported in Table V, which instead refers to the within-subject analysis with repeated measures ANOVA. Specifically, the p-values in Table V are the p-values of the F-statistic, which simply shows how likely is it that the null hypothesis of “no difference among group means” is true. Therefore, for this analysis, to the best of our knowledge, adjustment is not necessary. There is, however, a procedure for instead getting pairwise (e.g. isolated-baseline, social-baseline, social-isolated) adjusted p-values for within-subjects RMANOVA, but it is very complicated and will result in similar information to Fig. 5, which presents the pairwise comparisons across subjects, which is a more generalizable result than the within-subjects one.
@@ VI Discussion
Fig. 6 and Tables LABEL:table:cm1 - LABEL:table:cm4 suggest that for the average human participant, sharing indoor spaces with mobile quadrupedal robots was associated with statistically significant physiological changes indicative of acute stress, which provides a response to . Further, it was possible to distinguish, with higher than chance-level accuracy, specific characteristics (single v. multi robot and navigation v. search) about the human-robot encounters based on the ECG and EDA signals. These statistical correlations provide insight to research questions . These observations should be critically examined for potential insights for designers of robots and human-robot interaction.
For instance, we found that the average participant was more stressed by robots on navigation behavior than those on search behavior. One could hypothesize that this was due to robots on search inherently spending more time near the participants, giving them an opportunity to become more familiar with the robot. Indeed, familiarity, which is influenced by interaction duration and frequency, is one of the major factors behind perceived safety [akalin2023taxonomy]. However, to confirm this idea, further analysis of the data may be necessary (e.g. correlate human-robot distance with stress features to decouple navigation/search behavior from distance; interview and questionnaire data collected may contain information about how differently the subjects perceived robots on navigation vs. search). While less statistically significant, we also found that the average participant was more stressed encountering multiple robots travelling together rather than alone. This could suggest that whenever possible, mobile robots should travel alone, at least in similar physical settings as our experiment (i.e. closed indoors space occupied by humans), to maximize perceived safety. We also found that whether the subjects were alone or together was not associated with significant differences in physiological features of stress, responding to . This could indicate that the level of stress perecieved by people due to the co-location by mobile robots may not pose significant difference when they are deployed in environments where people work alone vs environments where multiple people work alongside each other.
@@
VII Conclusions
We presented results found from experimental human-robot encounters in an indoor apartment setting with members of the University of Texas at Austin. Particularly, we found that participants showed a measurable response to the robots over the baseline, multiple robots compared to a single robot, and under navigation compared with search behavior. Results were computed using statistical analysis of features found from electrodermal activity and heart rate variability recorded during experiments. One can hypothesize that the classifier may generalize to other HRI scenarios where participants are seated in a comfortable indoor environment when robots are present. Further, this classifier may provide a baseline from which to continue to develop an online classifier for acute mental stress, which may be used to control robot behavior. The complete open access database can be found at [dataset] and includes a plethora of rich time series data outside of the scope of this manuscript. For instance, one can manipulate the robot pose information to identify when the robots are in the line-of-sight of participants or the distance to participants. This data offer a new context from which to study proxemics. The dataset also includes audio intensity data from participant seating locations, which can further the investigation of the impacts of sound on physiological response. Ongoing work looks to combine the presented results with findings from interviews with participants in order to paint a more complete picture of human response to mobile robot encounters.
VIII Acknowledgments